OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 8 A. (−9, 1) B. (9, −1) C. (9, 1) D. (−1, 9)

jimthompson5910 (jim_thompson5910):

Hint: add the two equations to get x − y = 10 x + y = 8 ----------- 2x+0y = 18 so 2x = 18 ----> x = ?? once you've found x, use it to find y

OpenStudy (anonymous):

is it (9,1)?

OpenStudy (anonymous):

B. 9, -1

jimthompson5910 (jim_thompson5910):

you found that x = 9, that much is correct

jimthompson5910 (jim_thompson5910):

x + y = 8 9 + y = 8 .... .... .... y = ??

OpenStudy (anonymous):

oh so it's (9,-1), then?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

alrighty, thank you!